A space telescope in orbit around the Earth discovers a new planet, which is observed to move around the Sun by an angle of 4.72 milliradians in a year. Assuming a circular orbit, estimate the distance, in A.U., of the planet from the Sun




1
Expert's answer
2021-02-24T16:59:29-0500

"D=\frac{206265}{p}=\frac{206265}{4.72\cdot 10^{-3}}=43.7\cdot 10^6~\text{AU}."